OpenStudy (spicytaco1):

Will fan Which fractions are equivalent to -15/20 ? Choose all answers that are correct.(multiple choice) A. -2/3 B. 9/-12 C. -12/16 D. -25/30

OpenStudy (anonymous):

3/4 * 2/2 = 6/8 3/4 * 3/3 = 9/12 3/4 * 4/4 = 12/16 3/4 * 5/5 = 15/20 3/4 * 6/6 = 18/24 Therefore: 3/4 = 6/8 = 9/12 = 12/16 = 15/20 = 18/24

OpenStudy (anonymous):

15 and 20 both are divisible by 5 divide both of them by 5 to get answer